# Scalable Training of Graph Foundation Models for Atomistic Materials Modeling: A Case Study with HydraGNN

### v1

@@ IV-D Scalable HPO
The hyperparameter tuning has spanned important architectural hyperparameters described in Table III. The range of architectural hyperparameters covers regions of the hyperparameter space that allow to construct HydraGNN models of extremely diverse size, which include the SMALL model and the LARGE models described in Table II as extremes.
@@
TABLE III: Set of architectural HydraGNN hyperparameters tuned by scalable HPO.
@@
Figure 11 reports the validation mean absolute error (MAE) as a function of wall-clock time. The scattered distribution of blue dots (corresponding to values of the validation MAE for different HPO trials) shows that the HPO maintains a good degree of exploration throughout the entire execution. The red solid line indicates the minimum validation MAE obtained at a given time during the HPO run. The fact that the red line progressively lowers as time progresses confirms that new HPO trials progressively selected and evaluate help identify GFM architecture with better accuracy.
@@
Fig. 12 highlights dynamic HBM utilization by HydraGNN during a small HPO exercise utilizing 320 GPUs. This exercise ran for 2.5 hours and completed 64 total trials dynamically via co-scheduling of five trials at a time in parallel. The time-series traces shown in Fig. 12 document individual GPU memory usage on each assigned compute node with a dynamic high water mark peaking at 99% of available memory. The variability of memory utilization across different GPUs is due to the fact that different groups of GPUs (associated with different HPO trials) train HydraGNN models of different size, which affects the amount of GPU memory engaged at different stages of the model training.
Figure 12: GPU HBM memory consumption traces sampled via low-noise telemetry harness (per GCD on each assigned compute node) during HPO exercise executed on OLCF-Frontier.
IV-E Full training of best performing HydraGNN models identified by scalable HPO
The 10 best performing HydraGNN models identified by HPO have been fully trained for 40 epochs to reach convergence of the training. We report the trend of the training and testing loss function for all 10 models in Figure 13. The training loss functions flatten at the end of the training history, indicating that the models have reached their maximum predictive capacity. Moreover, the models seems to reach similar final accuracy, thereby confirming that the HPO has indeed thoroughly spanned the hyperparameter space and identified HydraGNN models with similar predictive performance.
@@ V Conclusion and Future Work
In this work we described our approach towards developing and training predictive GFMs by scaling the HydraGNN architecture on hundreds of millions of atomistic materials modeling data using two DOE leadership class supercomputers, viz. NERSC-Perlmutter and OLCF-Frontier. We discuss optimizations and tools used for developing a GFM and running hyperparameter optimization at large scale.
We use distributed data management capabilities to partition large volumes of data across distributed computing resources and efficiently exchange data samples across devices using low-latency communication methods. This helps preserve global data shuffling, which is crucial for maintaining good convergence of the GFM training. By scaling HPO on over 87% of the exascale OLCF-Frontier supercomputer, we have assessed the importance of thoroughly exploring a large set of hyperparameter configurations to identify HydraGNN architectures with high predictive accuracy. Moreover, access to exceptionally performing large scale computing facilities allowed us to develop and test ensemble UQ capabilities to measure the degree of confidence associated with the HydraGNN predictions. Performing HPO and ensemble UQ at unprecedented scale on supercomputing facilities confirms our computational readiness in using HydraGNN to develop trustworthy (i.e, accurate and confident) GFMs to support the US-DOE materials science needs by providing robust and transferable computational capabilities for AI-accelerated materials discovery and design.

### v2

@@ IV-D Scalable HPO
The hyperparameter tuning has spanned important architectural hyperparameters described in Table IV. The range of architectural hyperparameters covers regions of the hyperparameter space that allow to construct HydraGNN models of extremely diverse size, which include the SMALL model and the LARGE models described in Table III as extremes.
@@
TABLE IV: Set of architectural HydraGNN hyperparameters tuned by scalable HPO.
@@
Fig. 11 reports the validation mean absolute error (MAE) as a function of wall-clock time. The scattered distribution of blue dots (corresponding to values of the validation MAE for different HPO trials) shows that the HPO maintains a good degree of exploration throughout the entire execution. The red solid line indicates the minimum validation MAE obtained at a given time during the HPO run. The fact that the red line progressively lowers as time progresses confirms that new HPO trials progressively selected and evaluate help identify GFM architecture with better accuracy.
@@
To characterize the dynamic resource behavior of HPO, the user-space telemetry tool highlighted in Section IV-A was enabled to capture a variety of GPU metrics during a small HPO exercise utilizing 320 GPUs. Fig. 12 highlights the memory subsystem showing individual GPU memory usage on each assigned compute node with a dynamic high water mark peaking at 99% of available memory. The variability of memory utilization across different GPUs is due to the fact that different groups of GPUs (associated with different HPO trials) train HydraGNN models of different size, which affects the amount of GPU memory engaged at different stages of the model training.
Figure 12: GPU HBM memory consumption traces sampled via omnistat telemetry harness (per GCD on each assigned compute node) during HPO exercise executed on OLCF-Frontier.
IV-E Energy Profiling
To quantify the energy usage as a function of different model sizes, three training epochs of the SMALL, MEDIUM, and LARGE model configurations listed in Table III were completed with the omnistat telemetry tool sampling at one second intervals. Measurements consider the entire run duration including I/O for the initial data loading process. Each model executed using 1,024 GPUs on 128 compute nodes which is the minimum node count needed to accommodate memory requirements for the LARGE model configuration. The resulting GPU energy measurements as a function of model size are summarized in Table V. While total energy scales with the execution time, note that GPU utilization (occupancy) also influences the energy consumed. Table V includes mean utilization observed across all 1,024 GPUs. The LARGE case showed the highest GPU utilization–around 89%. The underling power histories used to compute total energy consumed for each model configuration are shown in Fig. 13. From these plots, we see evidence of the underlying training process with three epoch cycles visible in the power response. Furthermore, the increased GPU utilization for the larger models leads to increased GPU power demand with the LARGE model encountering peak power measurements in excess of 520 W (the peak TDP power for the AMD MI250 socket is 560W).
Model size Duration Mean GPU Utilization GPU Energy Consumed
SMALL 17 mins 12.5 % 14.0 kWh
MEDIUM 25 mins 46.0 % 42.7 kWh
LARGE 133 mins 88.9 % 366.6 kWh
TABLE V: Energy usage during training on OLCF-Frontier.
Figure 13: GPU Energy use over time for three models – SMALL (top), MEDIUM (middle), and LARGE (bottom). Each line represents one AMD MI250x.
IV-F Full training of best performing HydraGNN models identified by scalable HPO
The 10 best performing HydraGNN models identified by HPO have been fully trained for 40 epochs to reach convergence of the training. We report the trend of the training loss for all 10 models in Fig. 14. The training loss flattens at the end of the training history, indicating that the models have reached their maximum predictive capacity. The trained 10 models will be used to provide ensemble predictions in future tasks with uncertainty measurement.
@@ V Conclusion and Future Work
In this work we described our approach towards developing and training predictive GFMs by scaling the HydraGNN architecture on hundreds of millions of atomistic materials modeling data using two DOE leadership class supercomputers, viz. NERSC-Perlmutter and OLCF-Frontier. We discussed optimizations and tools used for developing a GFM and running hyperparameter optimization at large scale.
We used distributed data management capabilities to partition large volumes of data across distributed computing resources and efficiently exchange data samples across devices using low-latency communication methods. This helped preserve global data shuffling, which is crucial for maintaining good convergence of the GFM training. By scaling HPO on over 87% of the exascale OLCF-Frontier supercomputer, we have assessed the importance of thoroughly exploring a large set of hyperparameter configurations to identify HydraGNN architectures with high predictive accuracy. Moreover, access to exceptionally performing large scale computing facilities allowed us to develop and test ensemble UQ capabilities to measure the degree of confidence associated with the HydraGNN predictions. Performing HPO and ensemble UQ at unprecedented scale on supercomputing facilities confirms our computational readiness in using HydraGNN to develop trustworthy (i.e, accurate and confident) GFMs to support the US-DOE materials science needs by providing robust and transferable computational capabilities for AI-accelerated materials discovery and design.